OpenStudy (anonymous):

The height of a parallelogram is 2 units more than the base. The measure of the base of the parallelogram is 10 units. What is the area, in square units, of the parallelogram?

OpenStudy (anonymous):

@dpaInc @jim_thompson5910 you guys are the best please help :)

jimthompson5910 (jim_thompson5910):

The base is 10. So b = 10 The height is 2 units more than the base, so the height is 10+2 = 12, which means h = 12

jimthompson5910 (jim_thompson5910):

Now plug all this into the area formula A = bh A = 10*12 A = 120 So the area of the parallelogram is 120 square units.

OpenStudy (anonymous):

Gahh your so good :)

jimthompson5910 (jim_thompson5910):

You can think of the area formula for the parallelogram as the area formula for a rectangle (since a parallelogram is just a "slanted" rectangle...so to speak)

jimthompson5910 (jim_thompson5910):

why thank you

OpenStudy (anonymous):

Your welcome your honestly better then my teacher :)

jimthompson5910 (jim_thompson5910):

well that's both good and bad lol (hopefully your teacher isn't that bad though)

OpenStudy (anonymous):

she's okay.....but everytime i need help your my SAVIOR hehe

jimthompson5910 (jim_thompson5910):

lol well I'm glad I am